can someone help me by tuesday plss :(?
solve the problem.
Amy can clean the elephant enclosure in 4 hours, but it takes joel 6 hours. how long would it take them working together?

Answer by rwm(914)

x/4+x/6=1
x/2+x/3=2
3x+2x=12
5x=12
x=12/5=2.4 hours
check
2.4/4+2.4/6=1
.6+.4=1
ok